# footnotetext: *Equal Contributionfootnotetext: Corresponding authors

### v1

@@ 3.3 Stage-II: Enhancing Reasoning Quality
Although supervised fine-tuning can elicit reasoning behavior in models, it is often accompanied by an increase in hallucinations. Prior work lv2024coarse ; akbar2024hallumeasure attributes this phenomenon to a mismatch between the generation paradigms during training and inference. Specifically, models are trained to predict the next token conditioned on preceding ground-truth tokens, while at inference time they rely on their own previously generated outputs. This discrepancy becomes particularly pronounced in tasks requiring long-form reasoning. As illustrated in Figure 2, even a correct answer can emerge from a flawed reasoning process involving terminology misuse, logical errors, or fabricated medical content, etc. To align model outputs with human preference, the mainstream method is to employ Reinforcement Learning from Human Feedback (RLHF) ziegler2019fine ; ouyang2022training , exemplified by Preference optimization (PO). In this work, we introduce Multimodal Medical Reasoning Preference (MMRP) and integrate it with Direct Preference Optimization (DPO) rafailov2023direct ; xu2024dpo . This combination effectively reduces hallucinations and generates reasoning processes that align more closely with user requirements.
Preliminary Considering a model as a policy $\pi_{\theta}(y|x)$ parameterized by $\theta$, RLHF aims at aligning the LLM $\pi_{\theta}$ with human preference. DPO is a representative algorithm of RLHF, serving as the preference loss to optimize the policy $\pi_{\theta}$ by enabling the model to learn the relative preference between chosen and rejected responses. DPO eliminates the requirement of training an explicit reward model based on the assumption of the Bradley-Terry model huang2004generalized and directly optimizes $\pi_{\theta}$:
@@
Preference Data Construction. Initially, we prompt the Stage-I model $M_{1}$ to generate multiple candidate reasoning processes for each sample in our multimodal training dataset. Given an image $I$ and a question $q$, we sample candidate reasoning processes $y$ from the distribution $M_{1}(y\mid q,I)$, repeating this process 8 times per sample to form a candidate set $Y$. Notably, each question $q$ undergoes meticulous refinement through collaboration between human experts and a model (GPT-4o), explicitly removing descriptions of image $I$ from $q$. This ensures that reasoning arises from actual image examination rather than from textual captions. Subsequently, we employ GeminiPro2.5 to evaluate these candidate reasoning processes according to five MMRP criteria: Answer Correctness, Logical Consistency, Image Analysis Involvement, Absence of Hallucinations, and Presence of Reflection. Reasoning processes meeting all criteria form the positive set ($Y_{p}$), whereas those failing any criteria constitute the negative set ($Y_{n}$). We then construct preference pairs by selecting a preferred response $y_{c}$ from $Y_{p}$ and contrasting it with a response $y_{r}$ from $Y_{n}$. Unlike static dataset construction, our MMRP methodology enables dynamic and flexible curation of multimodal preference datasets, effectively balancing specialization and generalization in reasoning tasks. The prompt used for data evaluation with Gemini-2.5-Pro can be found in Appendix A.
After applying DPO-based fine-tuning to the constructed preference dataset, the model demonstrates improved alignment with desired reasoning patterns and shows enhanced performance across diverse multimodal clinical scenarios.
@@ 4.1 Experimental Setup
We select QwenVL2.5-7B and InternVL3.0-8B as the base models to evaluate the effectiveness of MedE2, resulting in two variants: (1) +Stage-I, which involves supervised fine-tuning on text-only data containing reasoning demonstrations; (2) +Stage-II, where the models from Stage-I are further enhanced using multimodal data and trained with DPO. We report the performance of GMAI-VL-R1 su2025gmai as a baseline method for comparison. For context, we report results from several leading proprietary models—GPT-4o, OpenAI-o1 openai2024reasoning , Gemini-2.5-Pro Google2025Gemini , and QvQ-Max teamqvq —to provide strong reference baselines. We also benchmark our approach against state-of-the-art open-source models that operate at substantially larger parameter scales bai2025qwen2 ; chen2024internvl .
All experiments are conducted on 8 NVIDIA H100 GPUs. LLaMA-Factory zheng2024llamafactory is used as the training framework, and vLLM kwon2023efficient serves as the backend during inference. In Stage-I, models are fine-tuned using LoRA hu2022lora with a learning rate of $1\times 10^{-4}$. Stage-II involves full-parameter fine-tuning with a learning rate of $1\times 10^{-6}$, optimized using DeepSpeed rasley2020deepspeed with the ZeRO-3 configuration. Both stages adopt a cosine learning rate decay schedule.
4.2 Evaluation Benchmarks
Our experiments involve two text-only tasks—MedQA jin2021disease and Medbullets chen2025benchmarking —as well as three multimodal benchmarks: MedXpertQA-MM zuo2025medxpertqa , MMMU-Health yue2024mmmu , and MMMU-Pro-Health yue2024mmmupro . Both MedQA and Medbullets are derived from questions used in the United States Medical Licensing Examination (USMLE). MedQA includes questions from both Step 1 and Step 2/3, while Medbullets focuses solely on Step 2/3, which are generally considered more challenging. For multimodal medical evaluation, MedXpertQA-MM serves as a highly challenging benchmark that assesses both medical knowledge and reasoning. MMMU-Health and MMMU-Pro-Health are healthcare-specific subsets of MMMU and MMMU-Pro, respectively.
4.3 Main Results
Text-only SFT Elicits Reasoning Behavior. As shown in Table 1, even a limited number of samples formatted with extended reasoning chains can effectively elicit the reasoning behavior, resulting in substantial performance gains. For example, QwenVL2.5-7B achieves improvements of +4.45% on MedXpertQA-MM and +6.67% on MMMU-Health. More impressively, the performance improvements from reasoning elicitation are even more pronounced for stronger base models. For example, InternVL3.0-8B achieves an 8% gain on MMMU-Health. This observation aligns with our hypothesis that training samples can serve as templates, illustrating how to utilize existing knowledge to solve complex reasoning tasks. Compared to GMAI-VL-R1 SFT, which relies on a 10K multimodal samples, our method achieves a 6 percentage point increase on MMMU-Health and a 3.37 percentage point gain on MMMU-Pro-Health only one-fifth of the data in a text-only format. This underscores that high-quality, task-targeted supervision is more critical than sheer data quantity for developing robust reasoning abilities.
Multimodal DPO Enhances Reasoning Quality. When Direct Preference Optimization (DPO) is applied after SFT in Stage-II, it builds upon the structured reasoning patterns established during the earlier stage and further enhances the model’s output quality. Compared to Group Relative Policy Optimization (GRPO)-based tuning shao2024deepseekmath (e.g., GMAI-VL-R1 RLT), our method achieves notably better results: 26.55% vs. 23.80% on MedXpertQA-MM, 66.00% vs. 57.33% on MMMU-Health, and 38.81% vs. 33.45% on MMMU-Pro-Health. A similar pattern is also observed in InternVL3.0-8B. These results suggest that merely forcing the model to reason over limited-solution-space tasks during training is suboptimal, often leading to hallucinations or incoherent reasoning. Instead, activating reasoning capabilities requires not only complex tasks but also training samples that engage the model in inference-time computation, or in other words, examples with precisely orchestrated solutions. Although the performance gain from DPO is slightly smaller than those from text-only SFT, we observe that DPO helps regulate the issue of “endless thinking” and encourages the model to “look before it thinks.” Qualitative analysis in Section 6 further shows that models trained with Stage-II generate more coherent and sophisticated reasoning processes.
@@ 4.3 Main Results
Method MedXpertQA-MM MMMU-Health MMMU-Pro-Health
@@ 4.4 Ablation Studies
Larger Models, Greater Improvements Based on the preceding results, we observe that models exhibiting stronger initial performance tend to benefit more from reasoning elicitation. We hypothesize that larger models equipped with more extensive pretrained knowledge can derive increased benefits from our proposed MedE2 framework. To rigorously verify this hypothesis, we apply our training pipeline to QwenVL2.5-32B and QwenVL2.5-72B and evaluate them on MedXpertQA-MM. Notably, these two larger models are fine-tuned using LoRA. We select MedXpertQA-MM for evaluation due to its increased challenge and minimal data leakage risk, as it was carefully constructed using difficulty-based filtering and data augmentation. Figure 3 illustrates the accuracy variations across different model sizes at various stages. Following Stage-I, QwenVL2.5-72B demonstrates the most substantial performance gain, outperforming both QwenVL2.5-32B (3.65%) and QwenVL2.5-7B (4.45%). These findings are consistent with our observations in Section 3.2: models with more extensive prerequisite knowledge encoded within their parameter space are more capable of learning reasoning patterns from limited yet precisely structured exemplars. Combining the results shown in Table 1, we surprisingly find that QwenVL2.5-72B enhanced with MedE2 even surpasses QvQ-Max (40.45% versus 36.25%) on the MedXpertQA-MM benchmark, which is currently one of the most powerful models within the Qwen series. Although larger proprietary models such as o1 and Gemini-2.5-Pro still maintain an advantage on challenging benchmarks, the application of MedE2 significantly narrows the performance gap between open-source and proprietary models. These results further demonstrate the effectiveness of MedE2 in advancing multimodal reasoning capabilities for medical tasks. Detailed experimental results can be found in Appendix B.
Figure 3: The performance of QwenVL2.5-7B, QwenVL2.5-32B, and QwenVL2.5-72B on the MedXpertQA-MM benchmark. As the model size increases, the models demonstrate progressively greater benefits from the MedE2.
@@
Figure 4: Comparison of performance on Med-XpertQA-MM using various strategies for eliciting reasoning behaviors.
@@
Benifition when Applying Inference Time Scaling. To investigate the effectiveness of MedE2, in scaling up inference time, we report the results of majority-vote accuracy and pass@n accuracy on MedXpertQA-MM and MMMU-Pro-Health across 32 computational runs with temperature =0.5. The results, illustrated in Figure 5, compare the performance of three models under Pass@n and majority voting (MV@n). Since both datasets are in a multiple-choice format with a limited answer space, model performance consistently improves as the number of attempts increases. Although the performance gain from majority voting on the MedXpertQA-MM dataset is modest due to its inherent difficulty, models trained with MedE2 still demonstrate a significant and stable advantage as the number of computational attempts increases. On MMMU-Pro-Health, employing majority voting yields notable performance enhancements. Interestingly, the accuracy of QwenVL2.5-7B initially rises to a peak but then slightly declines as computational runs increase. In contrast, models trained with MedE2 exhibit a steady performance improvement. These findings suggest that MedE2 enhances model output consistency, particularly evident during majority voting.
Figure 6: Case illustration of models’ reasoning capability in tackling clinical decision task.
@@ 4.5 Qualitative Analysis
Figure 6 demonstrates the reasoning capabilities of QwenVL2.5-72B+Stage2 through its coherent analysis of determining appropriate follow-up examinations based on the patient’s current clinical presentation. In real-world clinical scenatios, making an accurate diagnosis requires clinicians to integrate findings from multiple examinations to progressively narrow the differential diagnosis. Therefore, formulating an efficient and precise examination plan tailored to the patient’s chief complaint is critical. Taking a closer look to the case, a notable observation is that the reasoning process is not constrained by the given options. Instead, it reflects a stepwise, hypothesis-driven diagnostic approach rather than merely discussing each option sequentially. Specifically, the model first identifies potential diagnoses (e.g., Thrombocytopenia-Absent Radius syndrome and Fanconi anemia) and then systematically excludes one by evaluating evidence from the radiographic findings. Decisions regarding subsequent examinations are logically derived from the provisional diagnosis, aligning closely with realistic clinical workflows. This observed capability highlights the potential of MedE2 to serve as an effective post-trainig recipe to narrow the gap between current model performance and practical clinical application. We also remove the multiple-choice options and allow the model to respond freely, which can be seen in the Figure 7 of Appendix C.
5 Conclusion and Discussion
@@
Conclusion. In this study, we propose a novel reinforcement-learning-free training pipeline, MedE2, designed to progressively elicit and enhance multimodal reasoning capabilities in medical domains. MedE2 comprises two distinct stages: it first performs supervised fine-tuning on carefully curated text-only medical reasoning data, followed by preference-based optimization using multimodal data through Direct Preference Optimization. To support medical reasoning training, we constructed a dataset of 3.5K questions spanning both textual and multimodal clinical scenarios. Extensive experiments demonstrate the effectiveness and reliability of MedE2 in advancing the reasoning abilities of medical multimodal models. The demonstrated scalability and consistent improvements under inference-time scaling highlight its broad applicability across different model architectures and parameter scales. We hope that this early exploration into multimodal reasoning for medicine will inspire further research into specialized clinical reasoning capabilities.
Limitation and Impact. We hope that our post-training strategy can open new avenues for developing multimodal reasoning models in the medical domain. Despite the promising performance improvements demonstrated by our pipeline on current models, the lack of multimodal large language models specifically tailored to the medical domain presents significant challenges. Moreover, although current evaluation benchmarks are designed to resemble clinical scenarios, a notable gap remains. For example, many benchmark cases are crafted with sufficient cues that allow a medical specialist to deduce the correct diagnosis. In contrast, it would be challenging to create such a concise, complete and coherent case report in real clinical settings. Therefore, collecting real-world data and understanding how models generalize their reasoning to more complex and challenging clinical contexts remains a critical direction for future research.

### v3

@@ 3.3 Stage-II: Enhancing Reasoning Quality
Although supervised fine-tuning can elicit reasoning behavior in models, it is often accompanied by an increase in hallucinations. This poses significant risks in medical scenarios, where clinical decision-making requires strong evidence-based justification. Prior work (Lv et al., 2024; Akbar et al., 2024) attributes this phenomenon to a mismatch between the generation paradigms during training and inference. Specifically, models are trained to predict the next token conditioned on preceding ground-truth tokens, while at inference time they rely on their own previously generated outputs. This discrepancy becomes particularly pronounced in tasks requiring long-form reasoning. As illustrated in Figure 2, even a correct answer can emerge from a flawed reasoning process involving terminology misuse, logical errors, or fabricated medical content, etc. To align model outputs with human preference, the mainstream method is to employ Reinforcement Learning from Human Feedback (RLHF) (Ziegler et al., 2019; Ouyang et al., 2022), exemplified by Preference optimization (PO). In this work, we introduce Multimodal Medical Reasoning Preference (MMRP) and integrate it with Direct Preference Optimization (DPO) (Rafailov et al., 2023; Xu et al., 2024). This combination effectively reduces hallucinations and generates reasoning processes that align more closely with user requirements.
Preliminary Considering a model as a policy $\pi_{\theta}(y|x)$ parameterized by $\theta$, RLHF aims at aligning the LLM $\pi_{\theta}$ with human preference. DPO is a representative algorithm of RLHF, serving as the preference loss to optimize the policy $\pi_{\theta}$ by enabling the model to learn the relative preference between chosen and rejected responses. DPO eliminates the requirement of training an explicit reward model based on the assumption of the Bradley-Terry model (Huang et al., 2004) and directly optimizes $\pi_{\theta}$:
@@
Preference Data Construction. Initially, we prompt the Stage-I model $M_{1}$ to generate multiple candidate reasoning processes for each sample in our multimodal training dataset. Given an image $I$ and a question $q$, we sample candidate reasoning processes $y$ from the distribution $M_{1}(y\mid q,I)$, repeating this process 8 times per sample to form a candidate set $Y$. Notably, each question $q$ undergoes meticulous refinement through collaboration between human experts and a model (GPT-4o), explicitly removing descriptions of image $I$ from $q$. This ensures that reasoning arises from actual image examination rather than from textual captions. Subsequently, we filter out samples with wrong conclusions and employ Gemini-2.5-Pro to evaluate the remaining candidate reasoning processes according to four MMRP criteria: Logical Consistency, Image Analysis Involvement, Absence of Hallucinations, and Presence of Reflection. To validate the correctness of the model judgment, we select 1,000 cases and recruit human annotators to assess them using the same criteria. The differences between human scores and model judge scores are presented in Figure 3.
Figure 3: Distribution of humanmodel score differences, with 68.9% falling within $\pm\sigma$, where $\sigma$=1.02.
We find that model and human evaluations are largely consistent. Reasoning processes meeting all criteria form the positive set ($Y_{p}$), whereas those failing any criteria constitute the negative set ($Y_{n}$). We then construct preference pairs by selecting a preferred response $y_{c}$ from $Y_{p}$ and contrasting it with a response $y_{r}$ from $Y_{n}$. Unlike static dataset construction, our MMRP methodology enables dynamic and flexible curation of multimodal preference datasets, effectively balancing specialization and generalization in reasoning tasks. The prompt used for data evaluation with Gemini-2.5-Pro can be found in Appendix A. Consequently, we obtain 4,432 pairs for QwenVL2.5-7B and 5,224 for InternVL3.0-8B. After applying DPO-based fine-tuning to the constructed preference dataset, the model demonstrates improved alignment with desired reasoning patterns and shows enhanced performance across diverse multimodal clinical scenarios.
@@ 4.1 Experimental Setup
We select QwenVL2.5‑7B and InternVL3.0‑8B as the base models to evaluate the effectiveness of MedE2, resulting in two variants: (1) +Stage-I, which involves supervised fine-tuning on text-only data containing reasoning demonstrations; (2) +Stage-II, where the models from Stage-I are further enhanced using multimodal data and trained with DPO. We report the performance of GMAI‑VL‑R1 (Su et al., 2025) and Chiron-o1 (Sun et al., 2025) as baseline methods for comparison. We also benchmark our approach against state-of-the-art open-source models that operate at substantially larger parameter scales (Bai et al., 2025b; Chen et al., 2024b) and several leading proprietary models—GPT-4o, OpenAI-o1, Gemini-2.5-Pro, and QvQ-Max (Team, ). Details of Implementation can be found in Appendix C.
Evaluation Benchmarks. Our experiments involve two text-only tasks—MedQA (Jin et al., 2021) and Medbullets (Chen et al., 2025)—as well as three multimodal benchmarks: MedXpertQA-MM (Zuo et al., 2025), MMMU-Health (Yue et al., 2024a), and MMMU-Pro-Health (Yue et al., 2024b). Both MedQA and Medbullets are derived from questions used in the United States Medical Licensing Examination (USMLE). MedQA includes questions from both Step 1 and Step 2/3 of the Exam, while Medbullets focuses solely on Step 2/3, which are generally considered more challenging. For multimodal medical evaluation, MedXpertQA-MM serves as a highly challenging benchmark that assesses both medical knowledge and reasoning.
@@
Method MedXpertQA-$M$ MMMU-Health MMMU-Pro-Health
@@
4.2 Main Results
Text-only SFT Elicits Reasoning Behavior. As shown in Table 1, even a limited number of samples formatted with extended reasoning chains can effectively elicit the reasoning behavior, resulting in substantial performance gains. For example, QwenVL2.5-7B achieves improvements of +4.45% on MedXpertQA-MM and +6.67% on MMMU-Health. More impressively, the performance improvements from reasoning elicitation are even more pronounced for stronger base models. For example, InternVL3.0-8B achieves an 8% gain on MMMU-Health. This observation aligns with our hypothesis that training samples can serve as templates, illustrating how to utilize existing knowledge to solve complex reasoning tasks. Compared to GMAI-VL-R1 SFT, which relies on 10K multimodal samples, our method achieves a 6% increase on MMMU-Health and a 3.37% gain on MMMU-Pro-Health, using only half of the data in a text-only format. Similarly, relative to Chiron-o1, which relies on large-scale supervised fine-tuning with constructed CoT data, MedE2 exhibits a clear performance advantage across all benchmarks. This underscores that high-quality, task-targeted supervision is more critical than sheer data quantity for developing reasoning abilities.
Multimodal DPO Enhances Reasoning Quality. When Direct Preference Optimization (DPO) is applied after SFT in Stage-II, it builds upon the structured reasoning patterns established during the earlier stage and further enhances the model’s output quality. Compared to Group Relative Policy Optimization (GRPO)-based tuning (Shao et al., 2024) (e.g., GMAI-VL-R1 RLT), our method achieves notably better results: 26.55% vs. 23.80% on MedXpertQA-MM, 66.00% vs. 57.33% on MMMU-Health, and 38.81% vs. 33.45% on MMMU-Pro-Health. A similar pattern is also observed in InternVL3.0-8B. These results suggest that merely forcing the model to reason over limited-solution-space tasks during training is suboptimal, often leading to hallucinations or incoherent reasoning. Instead, activating reasoning capabilities requires not only complex tasks but also training samples that engage the model in inference-time computation, or in other words, examples with precisely orchestrated solutions. Although the performance gain from DPO is slightly smaller than those from text-only SFT, we observe that DPO helps regulate the issue of “endless thinking” and encourages the model to “look before it thinks.”
@@ 4.3 Ablation Studies
Larger Models, Greater Improvements Based on the preceding results, we observe that models exhibiting stronger initial performance tend to benefit more from reasoning elicitation. We hypothesize that larger models equipped with more extensive pretrained knowledge can derive increased benefits from our proposed MedE2 framework. To rigorously verify this hypothesis, we apply our training pipeline to QwenVL2.5-32B and QwenVL2.5-72B and evaluate them on MedXpertQA-MM. We select MedXpertQA-MM for evaluation due to its increased challenge and minimal data leakage risk, as it was carefully constructed using difficulty-based filtering and data augmentation. Figure 4 illustrates the accuracy variations across different model sizes at various stages. Following Stage-I, QwenVL2.5-72B demonstrates the most substantial performance gain, outperforming both QwenVL2.5-32B (4.65%) and QwenVL2.5-7B (4.45%). These findings are consistent with our assumption in Section 3.2: models with more extensive prerequisite knowledge encoded within their parameter space are more capable of learning reasoning patterns from limited yet precisely structured exemplars. Combining the results shown in Table 1, we surprisingly find that QwenVL2.5-72B enhanced with MedE2 even surpasses QvQ-Max (40.45% versus 36.25%) on the MedXpertQA-MM benchmark, which is currently one of the most powerful models within the Qwen series. Although larger proprietary models such as o1 and Gemini-2.5-Pro still maintain an advantage on challenging benchmarks, the application of MedE2 significantly narrows the performance gap between open-source and proprietary models. These results further demonstrate the effectiveness of MedE2 in advancing multimodal reasoning capabilities for medical tasks. Detailed experimental results can be found in Appendix D.
Figure 4: The performance of QwenVL2.5-7B, 32B and 72B on the MedXpertQA-MM benchmark. As the model size increases, the models demonstrate progressively greater benefits from the MedE2.
@@
Figure 5: Comparison of performance on Med-XpertQA-$M$ using various strategies for eliciting reasoning behaviors.
@@
Benefit when Applying Inference Time Scaling. To investigate the effectiveness of MedE2, in scaling up inference time, we report the results of majority-vote accuracy and pass@n accuracy on MedXpertQA-MM and MMMU-Pro-Health across 32 computational runs with temperature=0.5. The results, illustrated in Figure 6, compare the performance of three models under Pass@n and majority voting (MV@n). Since both datasets are in a multiple-choice format with a limited answer space, model performance consistently improves as the number of attempts increases. Although the performance gain from majority voting on the MedXpertQA-MM dataset is modest due to its inherent difficulty, models trained with MedE2 still demonstrate a significant and stable advantage as the number of computational attempts increases. On MMMU-Pro-Health, employing majority voting yields notable performance enhancements. Interestingly, the accuracy of QwenVL2.5-7B initially rises to a peak but then slightly declines as computational runs increase. In contrast, models trained with MedE2 exhibit a steady performance improvement. These findings suggest that MedE2 enhances model output consistency, particularly evident during majority voting.
Figure 7: Case illustration of models’ reasoning capability in tackling a clinical decision task.
@@ 4.4 Qualitative Analysis
Figure 7 demonstrates the reasoning capabilities of Our+Stage2 through its coherent analysis of determining appropriate follow-up examinations based on the patient’s clinical presentation. In real-world scenarios, making an accurate diagnosis requires clinicians to integrate findings from multiple examinations to progressively narrow the differential diagnosis. Taking a closer look at the case, a notable observation is that the reasoning process is not constrained by the given options. Instead, it reflects a stepwise, hypothesis-driven diagnostic approach rather than merely discussing each option sequentially. Specifically, the model first identifies potential diagnoses (e.g., Thrombocytopenia-Absent Radius syndrome and Fanconi anemia) and then systematically excludes one by evaluating evidence from the radiographic findings. Decisions regarding subsequent examinations are logically derived from the provisional diagnosis, aligning closely with realistic clinical workflows. This observed capability highlights the potential of MedE2 to serve as an effective post-training recipe to narrow the gap between current model performance and practical clinical application. Therefore, formulating an efficient and precise examination plan tailored to the patient’s chief complaint is critical. We also remove the multiple-choice options and allow the model to respond freely, which can be seen in the Figure 8 in Appendix E.
5 Conclusion
@@
In this study, we propose a novel reinforcement-learning-free training pipeline, MedE2, designed to progressively elicit and enhance multimodal reasoning capabilities in medical domains. MedE2 comprises two distinct stages: it first performs supervised fine-tuning on carefully curated text-only medical reasoning data, followed by preference-based optimization using multimodal data through Direct Preference Optimization. To support medical reasoning training, we constructed a dataset of 5K questions spanning both textual and multimodal clinical scenarios. Extensive experiments demonstrate the effectiveness and reliability of MedE2 in advancing the reasoning abilities of medical multimodal models. The demonstrated scalability and consistent improvements under inference-time scaling highlight its broad applicability across different model architectures and parameter scales. We hope that this early exploration into multimodal reasoning for medicine will inspire further research into specialized clinical reasoning capabilities.
Ethics Statement
This work complies fully with the ICLR Code of Ethics. All datasets used in this paper are publicly available and were obtained from established open sources. No private, sensitive, or personally identifiable information was collected or used. The study involves no human subjects, no experiments on vulnerable populations, and no interventions requiring IRB approval. We confirm that our methodology and results do not raise foreseeable risks of harm, misuse, or ethical concerns beyond standard scientific research practices.
Reproducibility Statement
We will open-source the curated dataset and model weights at https://anonymous.4open.science/r/MedEE-C381 to facilitate reproduction. We provide a comprehensive overview of our data construction methodology, the full processing pipeline, and the specific prompts used are provided in Appendix A. In addition, the training details and hyperparameter configurations for both stages of our method are presented in Appendix C.